# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

### v2

@@ 5.1. The Prediction Results
where Score is the difference between the AI agent’s predicted $Win\%$ and $Loss\%$. We calculated participants’ PredError for 17 predictions across all three games and their average PredError. A lower PredError indicates better prediction performance.
To compare participants’ prediction performance, we used mixed-design ANOVA for investigating both within-subject (to compare each participant’s 17 predictions) and between-subject (to compare All Original participants’ predictions vs. All Post-GenderMag participants’ predictions) factors. A mixed-design ANOVA can account for both the repeated measures within each participant and the differences between the two groups.
@@
Figure 12. The PredError for 17 prediction tasks across three games. G=Game, M=Move. For the six prediction tasks not faded—G1M9, G1M11, G2M6, G2M8, G3M7, and G3M13, Post-GenderMagand Originalparticipants’ PredError seemed to differ. Lower PredError indicates higher prediction accuracy.
@@ 5.2. Why the Difference? Explanations as both Friend and Foe
But, why did PredError flip-flop between the two groups for those six predictions? Explanations may hold the answer, because almost all differences between the two prototypes were fixes to explanations.
@@
(2) $\mu($ $\text{Mental Model Score}|\text{Explanation Usage})=\beta_{0}+\beta_{1}\times\text{Explanation Usage}$
where explanation usage is the frequency of each participant’s responses falling under the code “Participant used prior explanation” (Table LABEL:tab:codes_definitions).
Given that the explanations helped participants’ mental models, it seems logical that they would also help participants’ predictions. But what if, in some cases, participants followed the explanations “too closely”? For example, the “Top 5 moves” explanation not only highlights the importance of scores (which helped participants’ mental models), but also shows the moves that were the top 5 for the previous move. If the game remains on a similar trajectory as it was on that previous move, one of those top 5 is likely to be a good choice. However, if Agent Blue’s last move drastically changed the game’s situation, the previous moves will not be good predictions for the next one.
@@
What this shows us that the Post-GenderMag participants did what Team Game’s new explanations encouraged—basing their predictions on the top 5 moves. This strategy was often effective (the explanation was their “friend”) when the agent’s top moves from the previous move helped in predicting the future move.
For example, in Game 2 Move 6, Post-GenderMag group’s PredError tended to be lower than the Original group (Figure 12). This happened as Post-GenderMag participants selected a move “D3” (Figure 13) less frequently than the Original participants (3/35 Post-GenderMag participants vs 9/34 Original participants). This decision improved Post-GenderMag participants’ prediction accuracy since “D3” was a losing move for Move 6 and the correct prediction was “D4”. Post-GenderMag participants who followed the “top 5 moves” explanation too closely might have avoided “D3” as it was not in the top 5 moves listed (Figure 13). In contrast, Original participants, lacking this explanation, might have considered “D3” a good move based on the gameboard, as it creates multiple options for agent Blue and blocks agent Red.
Figure 13. Gameboard state and the top 5 moves that participants saw before before predicting Game 2: move 6. Here, the explanation “Top 5 Moves” was a friend for Post-GenderMag participants as they avoided selecting “D3” for not being in the top 5 list, which turned out to be a losing move next.
@@
But, when the top 5 moves did not list advantageous move, the explanation was “foe”. One such example was Game 1 Move 9. For this move, Post-GenderMag participants’ PredError was higher (Figure 12). This might have happened as Post-genderMag participants followed the top 5 moves more than the Original participants. 10/35 Post-GenderMag participants based their predictions on the top 5 moves compared to only 5/34 Original participants. This strategy did not work for the Post-GenderMag group, as “F3”, the correct prediction was outside the top 5. The top 5 moves might have misled the Post-GenderMag participants to overlook the gameboard potential of “F3”(Figure 14) so much that only 8/35 Post-GenderMag participants predicted “F3” compared 20/34 Original participants.
Thus, although the explanations were helpful for Post-GenderMag participants’ mental models, relying on them for predicting the future using the “Top 5 moves” did not always work.
6. RQ3 Results: Inclusivity
RQ3 asks whether the Post-GenderMag prototype provided better inclusivity support. So we investigated if the Post-GenderMag prototype affected participants with “Abi” and “Tim” facets (recall from Table 1) differently than the original version.
Although a facet-by-facet analysis produced raw numbers suggesting improvements for most facet values, no particular facet alone reached significance (Figure 15). However, when considering facet aggregates in individuals by gender, the inclusivity gains became clear.
We begin with two genders—women and men who made up about 93% of our study population, allowing us to perform inferential statistics. For participants from the LGBTQIA* community [^7] LGBTQIA* used based on Scheuerman et al.’s living document (Schmidt and Biessmann, 2020), we then provide a non-statistical representation of their data.
(a) Mental Model Score
@@ 6. RQ3 Results: Inclusivity
Figure 15. (a): Post-GenderMagAbi (stripes) and Tim (dots) participants had higher mental model scores than their Originalcounterparts. The five facets (from left to right): Info, Learn, Motiv, Risk, and SE (same order in adjacent plots).  (b): Post-GenderMagAbi’s and Tim’s PredError were higher but the difference from the Original group was not significant.
@@ 6.1. Inclusivity by Gender
Figure 16. Post-GenderMagparticipants’ mental model scores were higher than their Originalcounterparts, and the difference was significant for women. The X mark indicates the mean and the horizontal line indicates the median.
@@
Team Game considered Abi’s facets when making the fixes, so it’s not surprising that the fixes benefitted women as two-thirds ($67\%$) of the women had more Abi facets than Tim (Figure 17). In contrast, men tended more towards Tim ($57\%$) than Abi ($43\%$), yet Team Game’s fixes did not hurt men.
Figure 17. women(n=36) had more Abi facets/problem-solving styles whereas men(n=28) were more spread out but were slightly more Tim than Abi.
Now, what about equity? Equity is “the quality of being fair and impartial” (Press, 2019). Operationally, it would mean men and women achieving the same scores. Figure 18 illustrates the improvement in equity with the Post-GenderMag prototype. In the Original prototype, men and women had a difference of 44 points in their median mental model scores. This gap narrowed to just 24 points in the Post-GenderMag prototype—a 45% improvement in equity. So while improving inclusivity support, Team Game also made the Post-GenderMag prototype more equitable by gender.
Figure 18. Post-GenderMagmen and women show an improvement in equity compared to what Originalmen and women had. The X mark indicates the mean and the horizontal line indicates the median.
Other comparisons such as Post-Gendermag women’s PredError vs Original women’s PredError did not reveal any significant difference. Same thing happened for Post-Gendermag men vs Original men for PredError, mental model scores, and average explanation ratings.
For the comparisons, we used Student’s 2-sided t-tests when all three assumptions of independent samples t-test—1) independent observations, 2) normal distribution, and 3) equal variance were met. The independent observation condition was always true as each group was independent and the participants used the prototype individually with no interaction between them. When the normality distribution failed from a Shapiro-Wilk test, we conducted a non-parametric Wilcoxon Rank-Sum test to compare participants. Lastly, if equal variance failed from an F-test, we conducted a Welch’s t-test for the comparison.
Table 10 reports the GenderMag facet values of the five participants from the LGBTQIA* community. While the dataset is small, we hope it contributes to the growing literature on LGBTQIA* participants (e.g., (Acena and Freeman, 2021; Freeman et al., 2022; Hardy and Vargas, 2019)) and helps future researchers expand our understanding of inclusive design across all gender identities.
Table 10. LGBTQIA* facet values: Each row shows one LGBTQIA* participant’s GenderMag facet values. W = Woman, T = Transgender, NB = Non-Binary, PNS = Prefer Not to State.
PID W NB T PNS SE Motiv Learn Info Risk
P1 Higher Task by Tinkering Selective Tolerant
P11 Higher Tech by Process Selective Averse
P13 Lower Task by Process Comprehensive Tolerant
P46 Lower Task by Process Comprehensive Averse
P59 Higher Tech by Tinkering Comprehensive Tolerant
@@ 7. Discussion
7.1. Mental model scores improved but not prediction accuracy, why?
Our results showed that Team Game’s modified explanations helped improve participants’ mental models, but not their prediction accuracy. We wanted to understand why.
7.1.1. Heuristics vs Analytical thinking
The “dual-process theory of cognition” offers a useful lens for understanding the contradictory effects of XAI on our participants. This theory suggests that humans mostly rely on heuristics and mental shortcuts for everyday decisions, and rarely engage in analytical thinking due to its cognitive demand (Buçinca et al., 2021; Kahneman, 2011; Wason and Evans, 1974). Building on this theory, Buçinca et al. developed cognitive forcing functions for XAI to reduce AI overreliance (Buçinca et al., 2021). They posited that explanations primarily help users decide whether and when to follow AI recommendations based on general heuristics, rather than thinking analytically.
This might be true for our participants as well. What if Team Game’s modified explanations only supported participants’ heuristics and could not trigger deeper analytical thinking? We could not confirm this with our data, which could help explain why PostGenderMag participants did not excel in cognitively demanding tasks like predicting the AI agent’s next move.
So, similar to Buçinca et al, would Team Game have benefitted from integrating cognitive forcing functions into the explanations? Buçinca et al.’s study results offer limited optimism. While their cognitive forcing functions reduced AI overreliance, the functions could not improve AI-assisted human decisions (Buçinca et al., 2021).
7.1.2. Declarative vs Procedural knowledge
Post-GenderMag participants achieving significantly higher mental model scores yet no improvement in prediction performance could also stem from what they learn from Team Game’s fixes. Team Game provided declarative knowledge with their fixes, but no procedural knowledge. Declarative knowledge covers facts and information about the “what” of something and procedural knowledge addresses the steps and processes of “how” to do something (Anderson, 2005). In their explanations, Team Game provided “what” the AI considered for making decisions such as win%, loss%, draw%, and score%. However, they did not provide “how” the AI made those decisions (e.g. algorithmic steps, decision paths, or information flows). For Post-GenderMag participants, knowing the “what” led to higher mental model scores. But, they still needed the “how” to improve prediction accuracy.
7.1.3. More ways to fail than to succeed
Also, our participants had more ways to fail. For example, when predicting the first move on an empty gameboard, a participant had only a chance of 1 in 36 to make the correct prediction. We gave them partial credits when they predicted multiple squares and at least one was correct. Still, achieving higher prediction accuracy was difficult. This suggests Team Game’s fixes helped participants excel in basic tasks like understanding the AI agent’s behavior but not so much with advanced tasks like predicting the agent’s next move.
7.2. Abi vs Tim or both?—when making the fixes.
Team Game focused on Abi when finding and fixing the inclusivity bugs and it worked for women, who were predominantly Abi. Readers might wonder if Team Game also needed to consider Tim alongside Abi. We believe the answer is “‘no’ is acceptable but ‘yes’ is better”, and here is why.
Previous GenderMag studies have shown that Abi provides the strongest lens for inclusivity evaluation (Vorvoreanu et al., 2019; Guizani et al., 2022). These studies found more inclusivity bugs with Abi than with Tim, as Abi’s problem-solving styles often required additional inclusivity support. While considering Tim might have usually revealed a few additional inclusivity bugs, the additional time and effort could make the inclusivity evaluation process less feasible. Thus, we recommend following the “Abi-first” approach when the time is too limited to do both.

### v3

@@ 5.1. The Prediction Accuracy Results
where Score is the difference between the AI agent’s predicted $Win\%$ and $Loss\%$. We calculated participants’ PredError for 17 predictions across all three games and their average PredError. A lower PredError indicates better prediction accuracy.
To compare participants’ prediction accuracy, we used mixed-design ANOVA for investigating both within-subject (to use each of a participant’s 17 predictions as a data point) and between-subject (to compare the Original group’s predictions vs. the Post-GenderMag group’s predictions). A mixed-design ANOVA can account for both the repeated measures within each participant and the differences between the two groups.
@@
Details of the two groups’ PredError scores by task are shown in Figure 12. Note that the two groups had almost identical PredErrors in 11 of these tasks (faded colors), with visibly different box sizes/positions in only 6 tasks (not faded).
Figure 12. PredError for 17 prediction tasks across three games. G=Game, M=Move. For the 11 faded prediction tasks, participants in both the Post-GenderMaggroup and the Originalgroup made almost identical predictions. But this was not true for the 6 prediction tasks not faded—G1M9, G1M11, G2M6, G2M8, G3M7, and G3M13. (Lower is better, because low PredError indicates high prediction accuracy.)
@@ 5.2. Why the Difference? Explanations As Both Friend and Foe
(2) $\mu($ $\text{Mental Model Concepts Score}|\text{Explanation Usage})=\beta_{0}+\beta_{1}\times\text{Explanation Usage}$
where explanation usage is the frequency of each participant’s responses meeting the definition of code “Participant used prior explanation” (Appendix’s Table LABEL:tab:codes_definitions).
Given that the explanations helped participants’ mental model concepts scores, it seems logical that they would also help participants’ PredError scores. But what if, in some cases, participants followed explanations “too closely”?
Our participant data enabled considering this question for one of the explanation differences between the two prototypes—the addition of the “Top 5 Moves” explanation (Fix-7) into the Post-GenderMag prototype. Two aspects of the “Top 5 Moves” explanation are pertinent to our question of why participants’ mental model concepts scores were better in the presence of this explanation but their prediction accuracy scores were worse. The first aspect is that this explanation’s explicit inclusion of the Agent’s actual scores emphasizes them. The concept of the Agent emphasizing its scores, which was part of the mental model concepts scoring (recall Table 7), was one that was significantly more salient to the Post-GenderMag participants (as shown earlier in Figure 7). This evidence suggests that the “Top 5 Moves” explanation impacted Post-GenderMag participants’ mental model concepts scores positively.
The second pertinent aspect of “Top 5 Moves” is its emphasis on Agent Blue’s previous move’s top five. If a participant attended too much to that information, it could influence their prediction of Agent Blue’s next move either positively or negatively. It would be positive if the game remains on a similar trajectory as on the previous move, since those top 5 are likely to be good candidates for the next move. However, it would be negative if Agent Red’s move markedly changed the game’s situation, because the previous top 5 would not be likely to be good candidates in the present situation.
@@
What this shows is that the Post-GenderMag participants did exactly what Team Game’s Fix-7 explanation encouraged—basing their predictions on the Top 5 Moves explanation. This strategy was effective (the explanation was their “friend”) when the agent’s top 5 from the previous move helped in predicting the next move. This situation occurred in Figure 12’s G2M6 (Game 2 Move 6), G2M8, and G3M7.
For example, in the move before Game 2 Move 6, one of the top 5 moves had been moving to G2, which Post-GenderMag participants saw on display when they were predicting the agent’s upcoming Move 6. It turned out that Post-GenderMag participants were about 9% more likely than Original participants to predict move G2, which suggests that they were influenced by the Top 5 Moves explanation. And, because G2 remained a pretty good choice for Move 6 (still highly ranked by the agent), this choice helped Post-GenderMag participants’ prediction accuracy, as Figure 12 shows.
However, when the Top 5 Moves for the previous move did not include an advantageous choice for the upcoming move, the explanation acted as a “foe”. One such example was Game 1 Move 9 (G1M9 in Figure 12). For this move, Post-GenderMag participants’ PredError was higher/worse, perhaps because Post-GenderMag participants followed the Top 5 Moves more than the Original participants: 10/35 Post-GenderMag participants based their predictions on the Top 5 Moves compared to only 5/34 Original participants. But the correct choice for this move, F3, was not in Top 5 Moves display, because the game situation had changed too much. Thus, the Top 5 Moves explanation from the last move was misleading (Figure 13) so much that only 8/35 Post-GenderMag participants predicted “F3” compared to 20/34 Original participants.
Thus, although the fixed explanations were helpful for Post-GenderMag participants’ mental model concepts scores, using them to predict the next move sometimes worked against them. Our results can be seen as a nuanced case of overreliance—relying on the explanations to build conceptual understanding turned out to be appropriate, but relying on them to predict what the agent will do next turned out not to be generally appropriate.
6. RQ3 Results: Inclusivity
For curb-cut effects to be present, the effects need to be both “better for the underserved” population and “better for everyone else too.” The previous two sections investigated the latter effect, and found that, although the Post-GenderMag group did not achieve better prediction accuracy, the entire Post-GenderMag group did benefit with higher mental model conceptual understanding. This section considers the “underserved” effects, i.e., whether the underserved population also benefited with higher mental model conceptual understanding.
6.1. Inclusivity for Abi-Like Problem-Solvers
In this study, “Abi-like” users are the underserved population, because that is the population Team Game focused on supporting better with their inclusivity bug fixes. RQ3 asks whether the Post-GenderMag prototype provided better inclusivity for these particular users.
To investigate this question, we statistically tested mental model concept scores using the statistical reasoning explained in Section 3.3. We compared the mostly Abi-like subset of the Post-GenderMag group (participants with 3, 4, or 5 Abi-like problem-solving styles as per their questionnaire responses) vs. the mostly Abi-like subset of the Original group. We hypothesized that GenderMag-driven inclusivity fixes would best support the mostly Abi-like Post-GenderMag group’s mental model concepts scores.
@@ 6.1. Inclusivity for Abi-Like Problem-Solvers
Figure 14(a) shows a close-up view of these improvements, one problem-solving style at a time. At this level of detail, inferential statistics are not viable, but the raw numbers are still illustrative. As the figure shows, for participants with each of the Abi-like problem-solving values, the Post-GenderMag participants’ average mental model concepts scores were always higher than the Original participants’. That the Abi-like gains were visibly higher than the Tim-like gains for every one of the problem-solving style types confirms the “especially for underserved” aspect of curb-cut effects. That this was also true to a lesser extent for all but one of the Tim-like problem-solving values confirms the “for everyone else” aspect of these inclusivity fixes’ curb-cut effects.
On the other hand, as Figure 14(b) shows, prediction errors abounded with the new fixes, and with no advantageous effects for Abi-like participants or for anyone. This is unsurprising, given the results of Section 5. We will return to prediction errors in Section 7.
(a) Mental Model Concepts Score
@@
Figure 14. (a): Post-GenderMagAbi (striped) and Tim (dotted) participants had higher (better) mental model concepts scores than their Originalcounterparts. The five problem-solving styles (from left to right): Info, Learn, Motiv, Risk, and SE (same order in adjacent plots). (b): Post-GenderMagAbi’s and Tim’s PredError were higher (worse) but the difference from the Original group was not significant. x: mean, —: median.
@@ 6.2. Inclusivity by Gender
Investigating curb-cut effects through the lens of gender inclusivity requires identifying the underserved gender. Recall that the GenderMag inclusivity method has a dual focus: it aims to address problem-solving inclusivity in order to also address gender inclusivity. Thus, because Team Game used GenderMag’s Abi persona, which was designed to capture problem-solving styles more common among women than other people, Abi-like users being the underserved population also suggests women to be an underserved population.
We begin with the two genders who made up about 93% of our study population, thereby allowing inferential statistics—women and men. For participants in the LGBTQIA* community [^7] LGBTQIA* used based on Scheuerman et al.’s living document (Schmidt and Biessmann, 2020)., we then provide a non-statistical representation of their data.
6.2.1. Statistical Gender Results
@@
Figure 15. Post-GenderMagmen’s and women’s mental model concepts scores were higher than their Originalcounterparts. The difference was significant for women. x: mean, —: median.
@@
The relationship between problem-solving styles and gender also helps to explain the higher inclusivity for women. Team Game aimed their fixes toward including Abi’s problem-solving styles and, as Figure 16 shows, women were particularly likely to report Abi-like problem-solving styles in their questionnaire responses. Specifically, two-thirds ($67\%$) of the women had more Abi-like than Tim-like problem-solving styles (Figure 16). In contrast, men tended more towards Tim ($57\%$) than Abi ($43\%$). Thus, the fact that men’s scores also increased slightly is again evidence of a curb-cut effect.
Figure 16. 67% of women(n=36) had more Abi-like than Tim-like problem-solving styles, compared to 43% of men(n=28); and 33% of women had more Tim styles, compared to 57% of men.
A related concept to inclusivity is equity. Equity is “the quality of being fair and impartial” (Press, 2019). In this study, equity would mean men and women achieving the same scores. As Figure 17 shows, although Team Game’s inclusivity fixes did not fully achieve equity, they did improve it. With the Original prototype, women’s median mental model concepts scores were 44 points lower than men’s, but with the Post-GenderMag prototype, the gap narrowed to just 24 points—a 45% improvement in equity.
Figure 17. Post-GenderMagmen and women show an improvement in equity compared to what Originalmen and women had. x: mean, —: median.
6.2.2. LGBTQIA* Results
Table 9 reports the problem-solving style values and results of the five participants from the LGBTQIA* community. As with many other such reports of LGBTQIA* data (e.g., (Anderson et al., 2024)), the dataset is too small for inferential statistics. Still, we hope it contributes to the growing literature on LGBTQIA* participants (e.g., (Acena and Freeman, 2021; Freeman et al., 2022; Hardy and Vargas, 2019)) and helps future researchers expand the field’s understanding of inclusive design across all gender identities.
Table 9. LGBTQIA* problem-solving style values. Each row shows one LGBTQIA* participant’s GenderMag problem-solving style values. W = Woman, T = Transgender, NB = Non-Binary, PNS = Prefer Not to State, Select. = Selective, Comp. = Comprehensive, MMCS = Mental Model Concepts Score, PE = PredError.
PID W NB T PNS SE Motiv Learn Info Risk MMCS PE
P1 High Task Tinker Select. Tolerant 88 0.1224
P11 High Tech Process Select. Averse 207 0.142
P13 Low Task Process Comp. Tolerant 142 0.0557
P46 Low Task Process Comp. Averse 45 0.074
P59 High Tech Tinker Comp. Tolerant 174 0.1176
@@ 7. Discussion
7.1. When do predictions matter?
Prediction accuracy is a common metric for measuring users’ mental models of AI (Hoffman et al., 2019; Anderson et al., 2020), but we believe the use of this metric needs careful consideration. Predicting an AI agent’s next move requires a user’s mental model to be accurate enough to enable actually simulating an agent’s behavior. This raises the question of when this ability is needed.
Consider the example of self-driving cars. Potential buyers of such cars might need to be able to predict critical decisions like whether and when the car’s AI will hit the brakes if a pedestrian starts crossing the road in front of it. In contrast, buyers may feel no need to predict decisions like the AI’s route to reach some destination.
This suggests that evaluating users’ ability to predict an AI’s next action may be highly relevant for some XAI projects, but completely irrelevant in others. In our study, we directed participants to “understand” and “describe” the agent and to rate how well explanations helped them do these things. For this kind of explanation goal, predictions may not be a particularly relevant measure. On the other hand, suppose we had tasked participants to decide whether they would like the agent to “play future games on their behalf” (e.g., (Dodge et al., 2021a)). In that situation, users’ ability to simulate what an agent will do in certain situations may be extremely relevant as a measure of the approach’s effectiveness.
Thus, our call to the XAI research community is to rethink using predictions to measure a particular XAI project’s effectiveness—using predictions as a measure when the explanations actually aim to support prediction, but not using predictions to measure effectiveness when the project’s use-cases do not suggest the need for users to predict the AI.
7.2. Cognitive Lenses on Mental Model Concepts Scores vs. PredError Scores
We have just pointed out that there are situations in which explanations do need to support users’ ability to predict an AI’s next action. Two theories of cognition offer interpretations of why Team Game’s explanation fixes did not do so, even though they produced mental model improvements, and possible approaches to better enable explanations to support predicting.
The “dual-process theory of cognition” suggests that humans mostly rely on heuristics and mental shortcuts for everyday decisions, and rarely engage in analytical thinking due to its cognitive demand (Buçinca et al., 2021; Kahneman, 2011; Wason and Evans, 1974). Heuristic processing uses mental shortcuts such as simple decision rules to arrive quickly at a judgment, such as relying on currently held information or a perceived expert opinion (Trumbo, 1999). In contrast, analytical thinking (also called systematic thinking) involves effortful scrutiny and comparison of information, or as Trumbo puts it “carefully examining, comparing, and relating arguments…and a scrutiny of arguments, maintaining higher standards for the quality of information” (Trumbo, 1999). Building on this theory, Buçinca et al. posited that users tend to process explanations based on general heuristics rather than thinking analytically, which can lead to overreliance on AI decisions. For this reason, they developed cognitive forcing functions for XAI to force users into more analytical thinking and thereby reduce AI overreliance (Buçinca et al., 2021). Their empirical results showed that the cognitive forcing did reduce AI overreliance.
The distinction between heuristic vs. analytical thinking may apply to our results as well. We believe that in some cases, accurately predicting the AI’s next move would require analytical thinking, because it tends to require to comparing information from multiple AI decision points. It also seems likely that Team Game’s explanation fixes instead encouraged participants to rely even more on heuristic thinking because it displayed easily noticed “expert” information they may have perceived to be sufficient, in the form of the Top 5 Moves display.
An example from our study is Game 1 Move 9 (the most difficult to predict, according to Figure 12), in which the AI’s move turned out to be F3. Post-GenderMag participants were more than twice as likely as Original participants to pick a displayed Top 5 Move here, but F3 was not one of the Top 5 Moves on display (recall Figure 13), so heuristic processing of the Top 5 Moves explanation would not help. Rather, the information needed to accurately predict Game 1 Move 9 would require comparing the AI agent’s past four moves’ scores, win/loss/draw %s, and rankings, while also considering the gameboard state so as to block the opponent. This suggests the possibility of using cognitive forcing functionalities such as Buçinca et al.’s (Buçinca et al., 2021) to force users into analytical thinking, for XAI applications in which ability to predict an AI is important.
Another factor that may have been at play is the difference between “declarative knowledge” vs. “procedural knowledge.” Declarative knowledge covers facts about something (e.g., New York is north of the Miami), whereas procedural knowledge covers steps and processes of how to do something (e.g., to get to New York from Miami, first get on route…, then switch to…) (Anderson, 2005; Salaberry, 2018). Stated another way, declarative knowledge enables someone to describe something, whereas procedural knowledge enables them to actually do that thing (Salaberry, 2018). Thus, declarative knowledge is often classified as a “first stage” of learning that, with practice and experience, can eventually lead to the next stage of learning, procedural knowledge (Salaberry, 2018).
The mental model concepts scores were based on declarative knowledge—i.e., facts about the AI such as “Agent doesn’t consider opponent’s moves” (recall Table 8)—and the fact that the Post-GenderMag participants’ mental model concepts scores were higher suggests that Team Game’s explanation fixes were effective at promoting this declarative knowledge. The fact that Post-GenderMag participants’ PredError scores were not better despite their higher mental model concepts scores, could stem from a lack of procedural knowledge to be able to predict the next steps in a procedure (Rittle-Johnson et al., 2001; Braithwaite and Sprague, 2021). Education scholars have pointed out that procedural knowledge usually requires practice carrying out a task (Salaberry, 2018). This suggests that, for explanations to support users’ ability to actually simulate an AI, they need to be designed in ways that enable users to practice building exactly these skills. How to harvest education research’s approaches to creating this kind of user experience for applications where predictions matter is an open question.